OpenStudy (anonymous):

What is 33 and 1/3 % of 27

OpenStudy (eyust707):

.333 * 27

OpenStudy (eyust707):

"of" means times

OpenStudy (eyust707):

percent means "per one hundred"

OpenStudy (anonymous):

(100/3) * (1/100) * 27 = 9

OpenStudy (anonymous):

Is it right @Enmalou

OpenStudy (anonymous):

If I went to a shop and ther was 33 percent off a top that cost 27 pund how muck off would I get

OpenStudy (anonymous):

9 dollars

OpenStudy (anonymous):

No wait

OpenStudy (anonymous):

I don't understand how you do it

OpenStudy (anonymous):

What do you mean by 33 and 1/3 ??

OpenStudy (anonymous):

33 an one third percent

OpenStudy (anonymous):

Do you understand that 33 and 1/3 = 100/3 ?

OpenStudy (anonymous):

Yes but I don't understand how u times it

OpenStudy (anonymous):

ok We want to get the percentage of 100/3 then (100/3)*(1/100) = 1/3 right?

OpenStudy (anonymous):

@Enmalou

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok then what is the 1/3 of 27 ?? Is not 9 ?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@Enmalou

OpenStudy (anonymous):

Right :) but I still don't no how u did it

OpenStudy (anonymous):

MAN, let's make it (100*1)/(3*100) * 27 = 9

OpenStudy (anonymous):

GOT IT?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Is it help